Translate the following phrase into an algebraic expression using the variable m. Do not simplify. the cost of renting a car for one day and driving m miles if the rate is $44 per day plus 20 cents per mile
Answer #1

1 cent = 1/100 dollars

=> 20 cents = 20/100 = 0.2 dollars ( OR $0.2)

Given that, the cost of renting a car for one day and driving m miles if the rate is $44 per day plus 20 cents per mile.

Let y be Cost of renting a car.

Let D be a number of days and m be a number of miles.

Therefore, required expression is as follows:

y = 44 * D + 0.2 * m

=> y = 44 * 1 + 0.2 * m

=> y = 44 + 0.2m

Answer : Expression : y = 44 + 0.2m